How do I find the x-intercepts of the parabola with vertex (-3,16)and y-intercept (0,-20)
Answer by lwsshak3(6491)

How do I find the x-intercepts of the parabola with vertex (-3,16)and y-intercept (0,-20)
**
Standard equation for parabola: y=A(x-h)^2+k, with (h,k) being the (x,y) coordinates of the vertex.
..
Given equation:
y=A(x+3)^2+16
using given point (0,-20) to solve for A
-20=A(0+3)^2+16
9A=-36
A=-4
Equation of parabola:
y=-4(x-h)^2+k
Find x-intercepts
set y=0
0=-4(x+3)^2+16
4(x+3)^2=16
(x+3)^2=4
x+3±√4=±2
x=-5
or
x=-1
Ans:
x-intercepts: -5 and -1